# SAINT: Attention-Based Policies for Discrete Combinatorial Action Spaces

### v2

@@ 5.2 State-Dependent Sub-Action Dependencies
Table 1: Performance in CoNE as action dimensionality increases. SAINT consistently achieves the highest reward across all action space sizes. Factorized and autoregressive baselines plateau at substantially lower reward levels, while Wol-DDPG and A2C fail to learn viable policies.
Varying Dependence
To assess SAINT’s robustness to different levels of sub-action dependence, we incrementally increased the number of pits in the 12-dimensional CoNE environment. Higher pit densities impose stronger coordination requirements, as more sub-action combinations must be carefully selected to avoid pits. To ensure that a valid path from the start state to the goal always exists, pits were placed only in interior (non-boundary) states. We generate environments with 10%, 25%, 50%, 75%, and 100% of interior states occupied by pits. Note that even in the 100% setting, all boundary states remain pit-free, guaranteeing the existence of at least one (possibly inefficient) path to the goal region. We exclude A2C from this experiment due to the computational intractability of modeling such a large discrete action space (nearly 17 million actions) with a flat categorical distribution.
@@
5.3 Weak Sub-Action Dependencies with Complex Dynamics
To evaluate SAINT in environments where sub-action dependencies are relatively weak but the underlying dynamics are complex, we consider discretized variants of the HalfCheetah, Hopper, and Walker2D MuJoCo locomotion tasks (Towers et al., 2024). In these environments, each continuous joint control signal is discretized into 11 bins, yielding large, structured action spaces while retaining the rich temporal and physical dynamics of the original tasks. Although the discretized action spaces are combinatorially large, prior work (Beeson et al., 2024) suggests that the dependencies among sub-actions are relatively weak in these domains. This setting thus provides a useful test of SAINT’s generality, elucidating whether the architectural overhead of self-attention remains beneficial when sub-action dependencies are weak or whether simpler factorized policies suffice.
Given the limitations of PPO and Wol-DDPG identified in Sections 5.1 and 5.2, we restrict our comparison to the pure factorized and autoregressive baselines, using identical PPO implementations to ensure a controlled evaluation.
(a) HalfCheetah
(b) Hopper
(c) Walker2D
Figure 3: Performance in discretized MuJoCo environments. While results are similar across methods in HalfCheetah, SAINT outperforms factorized and autoregressive baselines in Hopper and Walker2D, demonstrating its ability to handle complex action spaces even when sub-action dependencies are relatively weak.
As shown in Figure 3, SAINT matches baseline performance in HalfCheetah and achieves faster learning and higher returns in Hopper and Walker2D. This demonstrates that even when sub-action dependencies are weak, learning a set-based representation of the action space gives an advantage over the rigid assumptions of factorization or a fixed autoregressive order.
@@ 5.4 Offline RL
Finally, we evaluate whether SAINT can be used effectively as a policy architecture in offline RL. Specifically, we use the medium-expert datasets from the discretized DM Control tasks cheetah run, finger spin, humanoid stand, quadruped walk, and dog trot introduced by Beeson et al. (2024). Across these environments, the number of sub-actions ranges from six to 39.
In this section, we report results using Batch-Constrained Q-learning (BCQ) (Fujimoto et al., 2019) as the offline RL objective. Additional offline evaluations with AWAC and IQL are provided in Appendix E. As in Section 5.3, we limit our comparison to the pure factorized and autoregressive baselines.
@@
Table 3: Mean $\pm$ std performance on offline DM Control tasks with BCQ variants.
@@ 5.5 Analyses and Ablations
To evaluate SAINT’s design choices and robustness, we conduct three analyses. First, we compare FiLM-based state conditioning to alternative mechanisms. Second, we assess the trade-off between representational power and computational cost, quantifying how modeling sub-action interactions affects sample efficiency. These evaluations are performed in the CityFlow-Irregular environment and in the 10-dimensional CoNE setting with pits occupying 25% of interior states. Finally, we test SAINT’s robustness to architectural hyperparameters using the CityFlow-Irregular environment.
@@
We compare SAINT’s FiLM-based state conditioning to four alternatives: (1) applying cross-attention to the state before self-attention, (2) applying cross-attention to the state after self-attention, (3) interleaving cross-attention and self-attention layers, and (4) appending the state as an additional token within the sub-action self-attention block. As shown in Appendix C, FiLM achieves higher final performance and more stable training in CityFlow, and performs at least as well as the alternatives in CoNE. These results indicate that while multiple conditioning mechanisms are effective, FiLM provides a consistent performance advantage and stabilizes training.
@@
However, in practical settings efficiency is better measured by wall-clock time to reach a target return. The “time to factored performance” metric shows that SAINT and SAINT-IP reach the factorized baseline’s asymptotic return in less than one-third of the time in CityFlow and about 30% faster in CoNE. This indicates that the added computation of explicit action representations is outweighed by faster learning — by modeling dependency structure, SAINT achieves stronger policies in less time.
@@
   Time to factored performance 395.76 $504.97$ $535.9$
Table 4: Total wall-clock training time and time to reach the factored baseline’s final performance, both measured in seconds. “Total training time” reflects the full duration of training, while “time to factored performance” measures how quickly each method reaches that baseline.
@@
We evaluated SAINT’s sensitivity to architectural hyperparameters by sweeping the number of self-attention blocks $\{1,3,5\}$ and attention heads $\{1,2,4,8\}$, for a total of twelve configurations. As shown in Appendix G, performance varied within a narrow range — the best setting (3 blocks $\times$ 1 head) outperformed the weakest (1 block $\times$ 8 heads) by only $\sim 7\%$. Even the weakest configuration exceeded all baselines, underscoring SAINT’s robustness to attention depth and head count.
@@ 6 Discussion and Conclusion
In this work we show that learning explicit representations of sub-action interactions is an effective and practical approach for control in combinatorial action spaces. SAINT represents actions as unordered sets and applies self-attention to capture sub-action dependencies, yielding expressive yet tractable policies. This modeling delivers substantial gains in sample efficiency, accelerating convergence to high-performing policies. In environments with up to 17 million joint actions, SAINT consistently outperforms baselines that assume independence, impose ordering, or learn flat policies, demonstrating the effectiveness of modeling sub-action interactions for scalable combinatorial control.
@@ A.1 Environmental Setup
In both CityFlow-Linear (Figure 4(a)) and CityFlow-Irregular (Figure 4(b)), the state is represented as a flat integer vector, in which each value indicates the number of waiting vehicles on an incoming lane and its paired outgoing lane. The reward at each step is the negative of the average “pressure” across intersections, where an intersection’s pressure is defined as the absolute difference between its total incoming and outgoing vehicle counts. Pressure is a standard metric in traffic signal control literature, commonly used to quantify imbalance in intersection flow Lioris et al. (2016); Varaiya (2013).
@@ A.2 Learning Curves Including Wol-DDPG
Figure 5 presents the full training curves for all methods in the CityFlow environments, including Wol-DDPG. As noted in Section 5.1, Wol-DDPG consistently underperforms relative to other methods. This poor performance is consistent with prior findings Chen et al. (2023), which show that Wol-DDPG is ill-suited to settings with unordered sub-actions. Learning curves excluding Wol-DDPG are presented in Figure 2.
@@ B.1 Environmental Setup
The Combinatorial Navigation Environment (CoNE) Landers et al. (2024) is designed to evaluate RL algorithms in settings with high-dimensional, combinatorial action spaces and strong, state-dependent sub-action dependencies. In CoNE, actions are formed by simultaneously selecting discrete sub-actions, each specifying movement along a different dimension. These sub-actions are executed in parallel to produce a composite transition, which may advance the agent toward the goal or result in failure by entering a pit.
@@
To our knowledge, no other existing benchmarks offer the combination of large-scale action spaces and tunable sub-action dependencies found in CoNE. Popular environments such as the DeepMind Control Suite Tassa et al. (2018), for example, lack meaningful sub-action interactions Beeson et al. (2024).
@@ B.3 Varying Dependence
Appendix C State Conditioning
(a) CityFlow
(b) 10-D CoNE with 25% pits
Appendix D Comparison to CTDE Architecture

### v3

@@ 5.1 State-Independent Sub-Action Dependencies
Table 1: Performance in CoNE as action dimensionality increases. SAINT achieves the highest reward across all action space sizes. Factorized and AR baselines plateau at substantially lower reward levels, while Wol-DDPG and A2C fail to learn viable policies.
To evaluate SAINT in environments where sub-action dependencies are primarily state-independent, we use the CityFlow traffic control benchmark (Zhang et al., 2019), where each action corresponds to simultaneous phase decisions across multiple intersections. While coordination is necessary for global traffic efficiency, the structure of these dependencies remains largely unchanged across states.
The results in Figure 2, demonstrate the advantage of SAINT’s set-based action representation. In both environments, SAINT learns faster and achieves higher final performance. The factorized approach is structurally incapable of modeling coordination between intersections and thus performs poorly. The autoregressive model is restricted by its fixed sequential prior, which is misaligned with the unordered nature of traffic-signal control. The flat PPO baseline, which cannot exploit combinatorial structure, fails to learn an effective policy.
5.2 State-Dependent Sub-Action Dependencies
In this setting, SAINT, the factorized baseline, and the AR baseline use identical A2C implementations to ensure a controlled comparison.
@@ 5.2 State-Dependent Sub-Action Dependencies
Varying Dimensionality
Varying Dependence
To assess SAINT’s robustness to different levels of sub-action dependence, we incrementally increased the number of pits in the 12-dimensional CoNE environment. Higher pit densities impose stronger coordination requirements, as more sub-action combinations must be carefully selected to avoid pits. To ensure that a valid path from the start state to the goal always exists, pits were placed only in interior (non-boundary) states. We generate environments with 10%, 25%, 50%, 75%, and 100% of interior states occupied by pits. Note that even in the 100% setting, all boundary states remain pit-free, guaranteeing the existence of at least one (possibly inefficient) path to the goal region. We exclude monolithic A2C from this experiment due to the computational intractability of modeling such a large discrete action space (nearly 17 million actions) with a flat categorical distribution.
@@ 5.3 Offline RL
Finally, we evaluate whether SAINT can be used effectively as a policy architecture in offline RL. Specifically, we use the medium-expert datasets from the discretized DM Control tasks cheetah run, finger spin, humanoid stand, quadruped walk, and dog trot introduced by Beeson et al. (2024). Across these environments, the number of sub-actions ranges from six to 38, yielding joint action spaces spanning several orders of magnitude, from hundreds to $3^{38}\approx 1.35\times 10^{18}$ possible actions.
In this section, we report results using Batch-Constrained Q-learning (BCQ) (Fujimoto et al., 2019) as the offline RL objective. Additional offline evaluations with AWAC and IQL are provided in Appendix E. Given the limitations of PPO and Wol-DDPG identified in Sections 5.1 and 5.2, we restrict our comparison to the pure factorized and autoregressive baselines, using identical PPO implementations to ensure a controlled evaluation.
@@
Table 3: SAINT outperforms all baselines across all offline DM Control tasks.
@@ 5.4 Analyses and Ablations
To evaluate SAINT’s design choices and robustness, we conduct three analyses. First, we compare FiLM-based state conditioning to alternative mechanisms. Second, we assess the trade-off between representational power and computational cost, quantifying how modeling sub-action interactions affects sample efficiency. These evaluations are performed in the CityFlow-Irregular environment and in the 10-dimensional CoNE setting with pits occupying 25% of interior states. Finally, we test SAINT’s robustness to architectural hyperparameters using CityFlow-Irregular.
@@
We compare SAINT’s FiLM-based state conditioning to four alternatives: (1) applying cross-attention to the state before self-attention, (2) applying cross-attention to the state after self-attention, (3) interleaving cross-attention and self-attention layers, and (4) appending the state as an additional token within the sub-action self-attention block. As shown in Appendix D, FiLM achieves higher final performance and more stable training in CityFlow, and performs at least as well as the alternatives in CoNE. These results indicate that while multiple conditioning mechanisms are effective, FiLM provides a consistent performance advantage and stabilizes training.
@@
However, in practical settings efficiency is better measured by wall-clock time to reach a target return. The “time to baseline” metric shows that SAINT and SAINT-IP reach the factorized baseline’s asymptotic return in less than one-third of the time in CityFlow and about 30% faster in CoNE. This indicates that the added computation of explicit action representations is outweighed by faster learning — by modeling dependency structure, SAINT achieves stronger policies in less time. Full learning curves are provided in Appendix F.
@@
     Time to baseline 395.76 $504.97$ $535.9$
Table 4: Total wall-clock training time and time to reach the factored baseline’s final performance (in seconds).
@@
We evaluated SAINT’s sensitivity to architectural hyperparameters by sweeping the number of self-attention blocks $\{1,3,5\}$ and attention heads $\{1,2,4,8\}$, for a total of 12 configurations. As shown in Appendix G, performance varied within a narrow range — the best setting (3 blocks $\times$ 1 head) outperformed the weakest (1 block $\times$ 8 heads) by only $\sim 7\%$. Even the weakest configuration exceeded all baselines, underscoring SAINT’s robustness to attention parameters.
@@ 6 Discussion and Conclusion
In this work we show that learning explicit representations of sub-action interactions is an effective and practical approach for control in combinatorial action spaces. SAINT represents actions as unordered sets and applies self-attention to capture sub-action dependencies, yielding expressive yet tractable policies. This modeling delivers substantial gains in sample efficiency, accelerating convergence to high-performing policies. In environments with up to $1.35\times 10^{18}$ joint actions, SAINT consistently outperforms baselines that assume independence, impose ordering, or learn flat policies, demonstrating the effectiveness of modeling sub-action interactions for scalable combinatorial control.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ A.1 Environmental Setup
In both CityFlow-Linear (Figure 3(a)) and CityFlow-Irregular (Figure 3(b)), the state is represented as a flat integer vector, in which each value indicates the number of waiting vehicles on an incoming lane and its paired outgoing lane. The reward at each step is the negative of the average “pressure” across intersections, where an intersection’s pressure is defined as the absolute difference between its total incoming and outgoing vehicle counts. Pressure is a standard metric in traffic signal control literature, commonly used to quantify imbalance in intersection flow (Lioris et al., 2016; Varaiya, 2013).
@@ A.2 Learning Curves Including Wol-DDPG
Figure 4 presents the full training curves for all methods in the CityFlow environments, including Wol-DDPG. As noted in Section 5.1, Wol-DDPG consistently underperforms relative to other methods. This poor performance is consistent with prior findings (Chen et al., 2023), which show that Wol-DDPG is ill-suited to settings with unordered sub-actions. Learning curves excluding Wol-DDPG are presented in Figure 2.
@@ B.1 Environmental Setup
The Combinatorial Navigation Environment (CoNE) (Landers et al., 2024) is designed to evaluate RL algorithms in settings with high-dimensional, combinatorial action spaces and strong, state-dependent sub-action dependencies. In CoNE, actions are formed by simultaneously selecting discrete sub-actions, each specifying movement along a different dimension. These sub-actions are executed in parallel to produce a composite transition, which may advance the agent toward the goal or result in failure by entering a pit.
@@
To our knowledge, no other existing benchmarks offer the combination of large-scale action spaces and tunable sub-action dependencies found in CoNE.
@@ B.3 Varying Dependence
Appendix C Comparison to CTDE Architecture
@@
Appendix D State Conditioning
(a) CityFlow
(b) 10-D CoNE with 25% pits